OpenStudy (anonymous):

Megan and Julia can finish apiece of work in 15 days. Megan can do the job herself in 20 days . If julie wanted to do the job alone how long would it take her

OpenStudy (amistre64):

what time of the month is it?

OpenStudy (anonymous):

lmao what

OpenStudy (amistre64):

:) nuthin

OpenStudy (amistre64):

1/M + 1/N = 1/15 do you see why?

OpenStudy (amistre64):

15 days; working together, in one day 1/15 is done

OpenStudy (anonymous):

yeah kinda

OpenStudy (amistre64):

we know M can do 1/20 in a day; working alone takes her 20 days, so 1/20 in a day

OpenStudy (amistre64):

1/20 + 1/N = 1/15 ; now we can solve for N

OpenStudy (amistre64):

Why i called julie N instead of J i cant tell; i was thinking Nancy lol

OpenStudy (anonymous):

I absolutely hate math, im just learning at this today

OpenStudy (amistre64):

oh well, shell just have to be Nancy, cant undo whats been done lol \[\frac{1}{N} = \frac{1}{15}-\frac{1}{20}\] \[\frac{1}{1} = N(\frac{1}{15}-\frac{1}{20})\] \[\frac{1}{(\frac{1}{15}-\frac{1}{20})} = N\]

OpenStudy (amistre64):

\[\frac{1}{(\frac{1}{15}-\frac{1}{20})} = N\] \[\frac{1}{(\frac{4}{60}-\frac{3}{60})} = N\] \[\frac{1}{\frac{1}{60}} = N\] \[\frac{1}{\frac{1}{60}}*\frac{\frac{60}{1}}{\frac{60}{1}} = N\] \[60 = N\] Nancy is slowww